OpenStudy (anonymous):

The sum of $1000 was invested for 4 years and the interest was compounded annually. If this investment amounts to $1463.33 after 4 years, what was the interest rate?

OpenStudy (anonymous):

\[1463.33=1000(1+\frac{ r }{ 1 })^{4}\]

OpenStudy (anonymous):

r=0.825, but when i go to check my answer i dont get 1463.33

OpenStudy (perl):

looks good so far

OpenStudy (perl):

1463.33/1000 = ( 1 + r)^4

OpenStudy (perl):

did you get that far?

OpenStudy (anonymous):

1.46333-1=r^4

OpenStudy (anonymous):

Fourth root 0.46333=r r=0.825035441

OpenStudy (anonymous):

When i go to check my answer i dont get the $1463.33? what am i doin wrong?

OpenStudy (anonymous):

@ganeshie8

OpenStudy (perl):

you made a mistake

OpenStudy (anonymous):

where?

OpenStudy (perl):

1463.33/1000 = ( 1 + r)^4 you cant subtract 1

OpenStudy (perl):

that is not valid step

OpenStudy (perl):

order of operations, (1+r)^4 , exponent first

OpenStudy (anonymous):

oooohhhhhh no wonderrrrr

OpenStudy (perl):

since what you have is 1463.33/1000 = ( 1 + r)*(1+r)*(1+r)*(1+r)

OpenStudy (perl):

what you can do is take the 4th root

OpenStudy (perl):

1463.33/1000 = (1+r)^4 (1463.33 /1000 ) ^(1/4) = ( 1 + r )

OpenStudy (anonymous):

alright thanks for finding my mistake